Singing Fish Fine Foods has a current annual cash dividend policy of ​$4.00. The price of the stock is set to yield a return of 14 %. What is the price of this stock if the dividend will be paid a.  for 11 ​years? b.  for 17 ​years? c.  for 40 ​years? d.  for 50 ​years? e.  for 100 ​years? f.  ​forever?
Answer #1

Price = Dividend × (1 – 1/(1 + r)n ) / r

a. Price = $4.00 × (1 – 1/(1.14)11 / 0.14 = $21.81

b. Price = $4.00 × (1 – 1/(1.14)17 / 0.14 = $25.49

c. Price = $4.00 × (1 – 1/(1.14)40 / 0.14 = $28.42

d. Price = $4.00 × (1 – 1/(1.14)50 / 0.14 = $28.53

e. Price = $4.00 × (1 – 1/(1.14)100 / 0.14 = $28.57

f. Price = $4.00 / 0.14 = $28.57